Jamie Lee Jackson, age 26, is in her last semester of college and is waiting for graduation day that is just around the corner! It is the time of year again when Jamie Lee must file her annual federal income taxes. Last year, she received an increase in salary from the bakery, which brought her gross monthly earnings to $2,800, and also opened up an IRA, to which she contributed $200 last year. Her savings accounts earn 2% interest per year, and she also had received an unexpected $1,750 gift from her great aunt, Jamie was also lucky enough last year to win a raffle prize of $2,500, most of which was deposited into her regular savings account after paying off her credit card balance Estimate Jamie Lees current tax liability by completing the form below. Use the 2016 values for standard deduction ($6,300) and personal exemption ($4,050). Each answer must have a value for the assignment to be complete. Enter O for any unused categories urrent Financial Situation Assets: Checking account Savings account (Interest earned last year) Emergency fund savings Monthly Expenses: $300 $2,600 Rent obligation $7,400 Utilities $160 $150 Food $4,400 Gas/maintenance $135 account (Interest earned last year) IRA balance Contribution made last year $100 $400 $200 Credit card payment Savings allocation: Regular savings (monthly) S0 $170 $3,500 Ralny day savings (monthly $50 ar Llabllltles: Student loan Credit card balance Interest pald last year) Income Gross monthly salary Entertalnment: Cake decorating class Movies with frlends $55 $11,300 $0 $80 $75 $2,800 Answer is not complete Estimated Income Tax Gross income (wages, salary, investment income, and other ordinary income) Less adjustments to income (see current tax regulations) Equals adjusted gross income Less standard deduction (use 2016 amounts) 200 S (200) $6,300

Need help figuring out the GROSS INCOME please
Answer #1

Answer

Calculation of gross income

Monthly expenses=

Rent obligation=300

Utilities=160

Food=155

Gas maintenance=135

credit card payment=0

Total =750

Total expense= monthly expense+ entertainment expenses

750+130=880

Income

salary=2800

gift=1750

Total interest earned last year= 250(150+100)=250/12=20 approx monthly expenses.

gross income= salary + interest+ gift=4570.